OpenStudy (anonymous):

When you write 0.0000001 as a power of 10: the base is 1 and the exponent is -7. the base is 1 and the exponent is -6. the base is 10 and the exponent is -7. the base is 10 and the exponent is -6.

OpenStudy (zzr0ck3r):

10^-1 =.1 10^-2 = .01 .....

OpenStudy (anonymous):

thanks

OpenStudy (zzr0ck3r):

which one is it?

OpenStudy (anonymous):

bottom

OpenStudy (anonymous):

hello brother..........

OpenStudy (zzr0ck3r):

-2 gives 1 zero -3 gives 2 zeros what gives 6 zeros?

OpenStudy (anonymous):

.......idk sorry iam not good at math u should no that

OpenStudy (zzr0ck3r):

-2 1 -3 2 -4 3 -5 4 ? 6

OpenStudy (anonymous):

7

OpenStudy (zzr0ck3r):

-7

OpenStudy (anonymous):

the base is 1 and the exponent is -7.

OpenStudy (zzr0ck3r):

basis is 10

OpenStudy (zzr0ck3r):

1^-7 = 1^8=1^878787877676=1

OpenStudy (anonymous):

thanks i will have to reveiw this again

OpenStudy (anonymous):

10–9 = 1.000000001 0.0000000001 0.000000001 0.0000001

OpenStudy (anonymous):

10^-9

OpenStudy (zzr0ck3r):

8 zeros

OpenStudy (zzr0ck3r):

look at the pattern

OpenStudy (anonymous):

0.000000001

OpenStudy (anonymous):

is that it

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

10–9 = 1.000000001 0.0000000001 0.000000001 0.0000001

OpenStudy (anonymous):



OpenStudy (anonymous):

8 zeros

OpenStudy (anonymous):

C

OpenStudy (anonymous):

0.000000001

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

To write 10^–7 in standard form, you would write a decimal number with: 1s in the first 7 decimal places. 1s in the first 6 decimal places and 0 in the 7th decimal place. 0s in the first 6 decimal places and 1 in the 7th decimal place. 0s in the first 7 decimal places and 1 in the 8th decimal place.

OpenStudy (anonymous):

0.0000001

OpenStudy (anonymous):

how many 0's is there

OpenStudy (anonymous):

after the decimal

OpenStudy (anonymous):

7

OpenStudy (anonymous):

1

OpenStudy (anonymous):

other side

OpenStudy (anonymous):

6

OpenStudy (anonymous):

which answer has 6 0's

OpenStudy (anonymous):

d

OpenStudy (anonymous):

no

OpenStudy (anonymous):

6 0's 1 in 7th spot

OpenStudy (anonymous):

is it abcd

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

To write 10^14 in standard form, you would use: a 10 and fourteen 0s. a 1 and fourteen 0s. fourteen 10s. a 14 and ten 10s.